Fran and Jenny run a child care service because they use Fran's house, Fran earns $150 more each month than Jenny. How much money (m) will Jenny earn in a month when the service takes in $2300?
a. equation
b. solution

Answer by mananth(16949)

jenny earns x
Fran earns x+150
Total 2300
x+x+150=2300
2x+150=2300
2x=2150
/2
x=1075
Jenny earns $1075
Fran earns 1075+150=$1225